OpenStudy (anonymous):

For my parabola assignment: what is the relationship between the x-intercepts and the x coordinate of the turning point? :) :) PLEASE HELP

OpenStudy (campbell_st):

the line of symmetry is in the middle of the x - intercepts... the turning point occurs on the line of symmetry.... e...g. y = x^2 + x - 6 has x intercepts at x = -3 & 2 line of symmetry is x = -b/2a so x = -1/2 midpoint of -3 and 2 = -1/2 Using the 1st derivative to find the stationary point.. y' = 2x + 1 stationary point at x = -1/2